# Tensor-based Graph Learning with Consistency and Specificity for Multi-view Clustering

### v2

@@ Abstract
In the context of multi-view clustering, graph learning is recognized as a crucial technique, which generally involves constructing an adaptive neighbor graph based on probabilistic neighbors, and then learning a consensus graph to for clustering. However, they are confronted with two limitations. Firstly, they often rely on Euclidean distance to measure similarity when constructing the adaptive neighbor graph, which proves inadequate in capturing the intrinsic structure among data points in practice. Secondly, most of these methods focus solely on consensus graph, ignoring unique information from each view. Although a few graph-based studies have considered using specific information as well, the modelling approach employed does not exclude the noise impact from the specific component. To this end, we propose a novel tensor-based multi-view graph learning framework that simultaneously considers consistency and specificity, while effectively eliminating the influence of noise. Specifically, we calculate similarity distance on the Stiefel manifold to preserve the intrinsic properties of data. By making an assumption that the learned neighbor graph of each view comprises a consistent part, a specific part, and a noise part, we formulate a new tensor-based target graph learning paradigm for noise-free graph fusion. Owing to the benefits of tensor singular value decomposition (t-SVD) in uncovering high-order correlations, this model is capable of achieving a complete understanding of the target graph. Furthermore, we derive an algorithm to address the optimization problem. Experiments on six datasets have demonstrated the superiority of our method. We have released the source code on https://github.com/lshi91/CSTGL-Code.
@@ I Introduction
The advent of the big data era has driven the prominence of multimedia data that originates from diverse sources [1]. For example, multimedia retrieval involves the integral utilization of color, textures, and edges [2], while video summarization requires the incorporation of multiple videos from distinct perspectives [3]. In the context of multimedia retrieval, each type of media data (color, textures, edges) can be considered as a different view. Consequently, multimedia data can be treated as a representative class of multi-view data, which has found extensive applications in multi-view clustering [4, 5].
@@
As a crucial unsupervised task, multi-view clustering overcomes the limitations of single-view clustering, particularly when the view is incomplete or the view data is noisy [6, 7]. In contrast to single-view clustering, a particular focus in multi-view clustering is on how to fuse data from multiple views to attain an enhanced data representation [8]. Multi-view Subspace Clustering (MVSC) aims to achieve a common subspace representation from multiple views, and then cluster a collection of data points based on the resulting low-dimensional subspaces [9, 10, 11, 12]. To exploit both the low-rank and sparse properties of data, Brbić et al. [13] jointly enforced the associated constraints on the affinity matrix. In [14], the authors further investigated the $l_{0}$ quasi-norm-based regularization to address the overpenalized problem caused by $l_{1}$-norm. Motivated by the underlying assumption that multiple views intrinsically originate from one shared latent representation, there has has been an increasing interest towards the latent representation multi-view methods [15, 16, 17]. Luo et al. proposed an effective subspace learning framework that considers both consistent and specific representations [18]. More progress on multi-view subspace clustering methods can be found in [19, 20, 21, 22, 23, 24].
Graph-based methods also hold an important standing in multi-view clustering. Nie et al. conducted several studies exploring the assignment of weights to different graphs for highlighting the significance of key views [25, 26]. Wang et al. [27] constructed a unified graph matrix which serves to fuse data graph matrices of all views. In [28], a parameter-free method that allows for structured graph learning was investigated. This method is also beneficial in reducing dimensionality. Liang et al. [29] formulated the consistent and inconsistent graph information in a unified optimization model. In [30], Kang at al. presented a graph learning framework that is able to preserve a complete structure of data. Li et al. proposed a consensus graph learning framework [31]. Wen et al. investigated the graph completion issue for incomplete views [32]. Further works related to graph-based methods are encouraged to refer [33, 34].
Recently, tensor-based methods have promoted the advancement of multi-view clustering towards the extraction of high-order data information [35]. Wu et al. [36] proposed an essential tensor learning method based on Markov chain for multi-view clustering. Considering that weakly supervised information can likely be acquired in real-world scenarios, Tang et al. [37] incorporated the weakly supervised information when formulating the regularization constraints. By incorporating tensor with graph learning, several tensor based graph learning methods can be discovered in [38, 39, 40, 41].
@@
Most of the aforementioned graph-based methods can be considered as drawing inspiration from [42, 26]. In essence, these methods are variations of learning a consensus graph by fusing adaptive neighbor graphs from different views. Despite a few graph-based studies, such as [29], have attempt to consider both consistency and specificity, they model the similarity matrix with a consistent part and a specific part, without any further meticulous consideration. Specifically, they overlook the impact of noise, which could result in the specific information learned being contaminated with noise. To be more clear, these methods have the following limitations:
@@
1) The adaptive neighbor graph is constructed using Euclidean distance, which probably fails to characterize the inherent properties of data.
@@
2) Most methods do not consider view-specific information when learning the consensus graph. Although a few works have made preliminary explorations in this direction, the modeling approach does not exclude the influence of noise from the specific component. For more details, readers are encouraged to refer to the motivation subsection in Section III.
@@
To address the drawbacks previously discussed, we propose a novel tensor-based graph learning framework, namely Tensor-based Graph Learning with Consistency and Specificity (CSTGL). We begin with learning the neighbor graphs of different views on the Stiefel manifold. The similarity between data points is measured using Stiefel distance. In contrast to Euclidean distance, Stiefel distance takes into account the orthogonal properties on the manifold [43], thereby providing a more accurate reflection of the underlying data structure. To facilitate the subsequent graph fusion, we make an assumption that the neighbor graph of each view can be explicitly modelled by a consistent component, a specific component, and a noise component. Then, we formulate a new tensor graph learning framework that takes full advantage of both consistent and specific information, while effectively mitigating the influence of noise. Our model can exploit the high-order correlations among graphs, thanks to tensor singular value decomposition (t-SVD) [44]. CSTGL can be treated as a noise-free graph fusion framework, and our main contributions include:
@@
  • We learn the neighbor graph for each view by measuring the similarity distance on the Stiefel manifold, which enables the exploitation of the data’s inherent structure properties.
@@
  • By assuming that the neighbor graph of each view is composed of a consistent component, a specific component, and a noise component, we construct a complete tensor graph fusion framework that explicitly incorporates both consistent and specific tensor graphs, while effectively minimizing the impact of noise.
@@
Experiments demonstrate that CSTGL outperforms some SOTA baselines. Moreover, ablation studies confirm the benefits of using Stiefel manifold distance for similarity measurement, as well as the improvements gained from the integration of consistent and specific tensor graphs.
@@ II Notations and Preliminaries
In the following, we list the notations used in this paper and introduce some fundamental definitions related to tensor. Moreover, we briefly review some relevant works.
@@ II-B Fundamental Concepts
To facilitate the understanding of tensor-related concepts, we introducing the following definitions [44]:
@@
Definition 3 (Orthogonal Tensor): For a orthogonal tensor $\mathcal{Q}\in\mathbb{R}^{n\times n\times n_{3}}$, it satisfies
@@ II-C Related Work
Before further proceeding, it is necessary to review two fundamental multi-view methods respectively corresponding to graph-based learning and tensor-based learning, namely Parameter-weighted Multiview Clustering (PwMC) [26] and t-SVD Multi-view Clustering (t-SVD-MSC) [35].
II-C1 PwMC
Suppose we need to partition $n$ samples into $c$ clusters, for given similarity matrices $\mathbf{A}^{(v)}\in\mathbb{R}^{n\times n}$ from different views, PwMC learns the target graph $\mathbf{S}$ by solving the following objective function
$\underset{\alpha^{(v)},\mathbf{S}}{\mathrm{min}}\sum^{m}_{v=1}\alpha^{(v)}\lVert\mathbf{S}-\mathbf{A}^{(v)}\rVert^{2}_{F}+\gamma\lVert\bm{\alpha}\rVert^{2}_{2}$ (7)
$s.t.\;\alpha^{(v)}\geq 0,\bm{\alpha}^{T}\mathbf{1}=1,s_{ij}\geq 0,\mathbf{s}_{i}\mathbf{1}=1,$
$\qquad rank(\mathbf{L}_{s})=n-c$
where $m$ denotes the number of views, $\bm{\alpha}=[\alpha^{(1)},\alpha^{(2)},\cdots,\alpha^{(m)}]$, $\gamma>0$, and $rank(\mathbf{L}_{s})$ represents the rank of the Laplacian matrix $\mathbf{L}_{s}$. The Laplacian matrix is defined as $\mathbf{L}_{s}=\mathbf{D}_{s}-(\mathbf{S}^{T}+\mathbf{S})/2$, where $\mathbf{D}_{s}\in\mathbb{R}^{n\times n}$ denotes the degree matrix, which is diagonal, with the $i$-th diagonal element determined by $\sum_{j}(s_{ij}+s_{ji})/2$. Inspired by the idea of learning the target graph in PwMC, many variants have been explored [45, 46, 31].
II-C2 t-SVD-MSC
To capture the low rank tensor subspace and high-order correlations among views, t-SVD-MSC performs a rotation on the tensor and imposes a low-rank tensor constraint on the rotated tensor. Given a set of multi-view data matrices $\{\mathbf{X}^{(v)}\}^{m}_{v=1}$, with $\mathbf{X}^{(v)}\in\mathbb{R}^{d_{v}\times n}$ denoting the $v$-th view data with $d_{v}$ dimensions, the objective function of t-SVD-MSC is defined as:
$\underset{\mathbf{Z}^{(v)},\mathbf{E}^{(v)}}{\mathrm{min}}\lambda\lVert\mathbf{E}\rVert_{2,1}+\lVert\mathcal{Z}\rVert_{\circledast},$ (8)
$s.t.\;\mathbf{X}^{(v)}=\mathbf{X}^{(v)}\mathbf{Z}^{(v)}+\mathbf{E}^{(v)},v=1,\cdots,V,$
$\qquad\mathcal{Z}=\varphi(\mathbf{Z}^{(1)},\mathbf{Z}^{(2)},\cdots,\mathbf{Z}^{(V)}),$
$\qquad\mathbf{E}=[\mathbf{E}^{(1)};\mathbf{E}^{(2)};\cdots;\mathbf{E}^{(V)}],$
where $\mathbf{Z}^{(v)}\in\mathbb{R}^{n\times n}$ stands for the $v$-th view representation matrix, $\mathbf{E}^{(v)}$ is the $v$-th view error matrix, the function $\varphi(\cdot)$ serves to span a 3-mode tensor by stacking various representation matrix, with its dimensionality rotated to $n\times m\times n$.
Owing to the superior properties of t-SVD, t-SVD-MSC is capable of exploiting the comprehensive information among all the views. Motivated by the t-SVD based tensor multi-rank constraint, several multi-view tensor learning methods have been investigated [36, 47, 38, 48].
@@ III-A Motivation
Observation 1: In order to reflect the similarity among data samples, a common strategy is to learn the adaptive neighbor graph for $\mathbf{A}^{(v)}$ [31, 39]. However, most methods construct the adaptive neighbor graph using Euclidean distance, which may not effectively uncover the inherent structure of complex data.
Observation 2: We also notice that the majority of graph-based multi-view methods follow the learning scheme in (7) to attain a target graph from multiple views. Essentially, these methods focus on constructing a consensus graph by minimizing the discrepancy between the similarity graph and the target graph. This learning strategy notably emphasizes the importance of consistency among views. However, from another perspective, it seems to neglect the diversity inherent in different views.
Observation 3: Despite a few graph-based works, such as [29], have attempt to consider both consistency and specificity, they model the similarity matrix $\mathbf{A}^{(v)}$ with a consistent part $\mathbf{S}$ and a specific part $\mathbf{S}^{(v)}$, i.e., $\mathbf{A}^{(v)}=\mathbf{S}+\mathbf{S}^{(v)}$. However, this model exists a significant limitation. It does not take into account the impact of noise, which leads to the noise being assigned to the specific component $\mathbf{S}^{(v)}$ during the learning process. As a consequence, the specific information learned is impure. Ideally, the specific component should reflect the unique information of each view that contributes to achieve comprehensive information, rather than containing noise. Therefore, it is necessary to explicitly distinguish noise from specific information when modeling.
@@ III-B Model Formulation
Stiefel manifold: The Stiefel manifold represents the set of matrices with orthonormal columns. The Stiefel manifold is commonly used in optimization and geometry, particular in problems involving rotations and orthogonal transformations. Given two $d\times n$ orthonormal matrices $\mathbf{X}$ and $\mathbf{Y}$, the $l_{1}$-norm geodesic distance between them is defined as [49]
$d_{s}(\mathbf{X},\mathbf{Y})=d-\mathrm{tr}(\mathbf{X}^{T}\mathbf{Y}).$ (9)
Inspired from (9), a natural variant that involves two vectors $\mathbf{x}$ and $\mathbf{y}$ can be formulated to serve as a distance measure on the Stiefel manifold
@@
where $d$ can be interpreted as a threshold to guarantee the distance to be positive.
@@
From the above discussion, we can define a new formulation to learn the neighbor graph $\mathbf{A}^{(v)}$, that is
@@
where $\alpha$ denotes the positive balance parameter, and $\mathbf{a}^{(v)}_{i}\in\mathbb{R}^{n\times 1}$ is a column vector with its $j$-th element being $a^{(v)}_{ij}$. The main difference between the traditional neighbor graph learning and our method exists in the distance measure. In contrast to the traditional approach that utilizes Euclidean distance $\lVert\mathbf{x}^{(v)}_{i}-\mathbf{x}^{(v)}_{j}\rVert^{2}_{2}$ to measure similarity, our method uses Stiefel manifold distance, which more precisely reflects the intrinsic structure for real-world data.
After learning the neighbor graphs of all the views, we need to fuse these neighbor graphs. As mentioned in Observations 2 and 3 of the motivation subsection, most graph-based multi-view methods follows (LABEL:eq07) to lean a consensus graph, thereby ignoring the specific information of each view. While a few works consider both consistency and specificity, they fail to distinguish noise from specific information. To address these limitations, we suppose that each neighor graph $\mathbf{A}^{(v)}$ can be modelled by a consistent component, a view-specific component, and a noise component, i.e.,
@@
Remark 2: In contrast to the model presented in Observation 3, our model explicitly separates the noise term from the specific component, which guarantees that the specific information learned is useful and free from noise. In essence, our modelling technique contributes to noise-free graph fusion.
For capturing the high-order correlation among different neighbor graphs, we formulate the following tensor-based graph fusion framework:
@@
Remark 3: We impose $l_{2,1}$-norm on the noise tensor $\mathcal{E}$ to encourage the columns of each frontal slice $\mathbf{E}^{v}$ to be zero, which is expected to minimize the impact of noise [9]. We enforce the consistent tensor $\mathcal{S}_{1}$ with the $t$-SVD based tensor nuclear norm to exploit more shared high-order information across various neighbor graphs [35]. Moreover, we impose $F$-norm on the specific tensor $\mathcal{S}_{2}$ in order to ensure the preservation of view-specific information.

### v3

@@ Abstract
In the context of multi-view clustering, graph learning is recognized as a crucial technique, which generally involves constructing an adaptive neighbor graph based on probabilistic neighbors, and then learning a consensus graph for clustering. However, it is worth noting that these graph learning methods encounter two significant limitations. Firstly, they often rely on Euclidean distance to measure similarity when constructing the adaptive neighbor graph, which proves inadequate in capturing the intrinsic structure among data points in practice, particularly for high-dimensional data. Secondly, most of these methods focus solely on consensus graph, ignoring unique information from each view. Although a few graph-based studies have considered using specific information as well, the modelling approach employed does not exclude the noise impact from the common or specific components. To this end, we propose a novel tensor-based multi-view graph learning framework that simultaneously considers consistency and specificity, while effectively eliminating the influence of noise. Specifically, we calculate similarity using pseudo-Stiefel manifold distance to preserve the intrinsic properties of data. By making an assumption that the learned neighbor graph of each view comprises a consistent part, a specific part, and a noise part, we formulate a new tensor-based target graph learning paradigm for noise-free graph fusion. Owing to the benefits of tensor singular value decomposition (t-SVD) in uncovering high-order correlations, this model is capable of achieving a comprehensive understanding of the target graph. Furthermore, we derive an algorithm to address the optimization problem. Experiments on six datasets have demonstrated the superiority of our method. We have released the source code on https://github.com/lshi91/CSTGL-Code.
@@ I Introduction
The advent of the big data era has driven the prominence of multimedia data that originates from diverse sources [1, 2]. For example, multimedia retrieval involves the integral utilization of color, textures, and edges [3], while video summarization requires the incorporation of multiple videos from distinct perspectives [4]. In the context of multimedia retrieval, each type of media data (color, textures, edges) can be considered as a different view. Consequently, multimedia data can be treated as a representative class of multi-view data, which has found extensive applications in multi-view clustering [5, 6].
@@
As a crucial unsupervised task, multi-view clustering overcomes the limitations of single-view clustering, particularly when the view is incomplete or the view data is noisy [7, 8]. In contrast to single-view clustering, a particular focus in multi-view clustering is on how to fuse data from multiple views to attain an enhanced data representation [9, 10]. Multi-view Subspace Clustering (MVSC) aims to achieve a common subspace representation from multiple views, and then cluster a collection of data points based on the resulting low-dimensional subspaces [11, 12, 13, 14]. Instead of directly learning the joint representation from original features, latent multi-view methods first extract the latent representation from multi-view data, and then learn the subspace self-representation matrix [15]. These methods effectively address the issue of suboptimal representation in multi-view subspace clustering when data is insufficient [16, 17]. In addition, tensor-based methods have promoted the advancement of multi-view subspace clustering towards the extraction of high-order data information [18, 19, 20].
@@
In addition, graph-based methods also occupy an important position in multi-view clustering. Basically, they conform to the paradigm of constructing neighbor graphs for each view from the original data points [21, 22], followed by the fusion of these graphs [23, 24]. Most graph-based methods can be considered as drawing inspiration from [25]. In essence, these methods are variations of learning a consensus graph by fusing adaptive neighbor graphs from different views. Despite a few graph-based studies, such as [24], have attempted to consider both consistency and specificity, they model the similarity matrix with a consistent part and a specific part, without any further consideration. Specifically, they overlook the impact of noise, which could result in the common and specific information learned being contaminated with noise. To be more clear, these methods have the following limitations:
@@
1) The adaptive neighbor graph is constructed using Euclidean distance, which probably fails to characterize the inherent properties of data, especially for high-dimensional data that is commonly encountered in real world scenarios.
@@
2) Most methods do not consider view-specific information when learning the consensus graph. Although a few works have made preliminary explorations in this direction, the modeling approach does not exclude the influence of noise from both the common and specific components. For more details, readers are encouraged to refer to the motivation subsection in Section III.
@@
To address the drawbacks previously discussed, we propose a novel tensor-based graph learning framework, namely Tensor-based Graph Learning with Consistency and Specificity (CSTGL). We begin with learning the neighbor graphs of different views using the pseudo-Stiefel manifold distance. In contrast to Euclidean distance, the newly adopted distance measure is a simplified version of the true Stiefel manifold distance [26], which is capable of providing a more accurate reflection of the underlying data structure. To facilitate the subsequent graph fusion, we make an assumption that the neighbor graph of each view can be explicitly modelled by a consistent component, a specific component, and an error component. Then, we formulate a new tensor graph learning framework that takes full advantage of both consistent and specific information, while effectively mitigating the influence of noise. Our model can exploit the high-order correlations among graphs, thanks to tensor singular value decomposition (t-SVD) [27]. CSTGL can be treated as a noise-free graph fusion framework, and our main contributions include:
@@
  • We learn the neighbor graph for each view by using the pseudo-Stiefel manifold distance, which allows for a more effective exploitation of the underlying structure compared to Euclidean distance.
@@
  • By assuming that the neighbor graph of each view is composed of a consistent component, a specific component, and an error component, we construct a complete tensor graph fusion framework that explicitly incorporates both consistent and specific tensor graphs, while effectively minimizing the impact of noise.
@@
Experiments demonstrate that CSTGL outperforms some SOTA baselines. Moreover, ablation studies confirm the benefits of using the pseudo-Stiefel manifold distance for similarity measurement, as well as the improvements gained from the integration of consistent and specific tensor graphs.
@@ II Notations and Preliminaries
In the following, we list the notations used in this paper and introduce some fundamental definitions related to tensor. Moreover, we briefly review some relevant work on subspace-based and graph-based methods.
@@ II-B Fundamental Concepts
To facilitate the understanding of tensor-related concepts, we introduce the following definitions [27]:
@@
Definition 3 (Orthogonal Tensor): For an orthogonal tensor $\mathcal{Q}\in\mathbb{R}^{n\times n\times n_{3}}$, it satisfies
@@ II-C Related Work
II-C1 Subspace-based Methods
Brbić et al. jointly enforced the associated constraints on the affinity matrix to exploit both the low-rank and sparse properties of data [28]. The authors further investigated the $l_{0}$ quasi-norm-based regularization to address the overpenalized problem caused by $l_{1}$-norm [29]. Motivated by the underlying assumption that multiple views intrinsically originate from one shared latent representation, there has been an increasing interest towards the latent representation multi-view methods [16, 30, 17]. In addition, some studies emphasize both consistent and specific representations to fully leverage the information from multi-view data [31, 32]. Considering that weakly supervised information can likely be acquired in real-world scenarios, Tang et al. incorporated the weakly supervised information when formulating the regularization constraints [33]. Applying kernel methods to subspace clustering enhances the data fusion from multiple views in nonlinear subspaces [34, 35]. In tensor-based multi-view methods, imposing low-rank constraints on tensors helps reduce redundancy among different views [36, 37, 38]. This is achieved by capturing essential information and discarding irrelevant or redundant features [38, 39]. Some researchers have also explored low-rank tensor approximation to enhance learning effectiveness in the presence of noise [40, 41, 42]. Further progress in multi-view subspace clustering methods can be found in the preservation of structural information [43, 44, 45], robustness inspired by information-theoretic principle [46, 47, 48], and mining of nonlinear data properties [49, 50].
II-C2 Graph-based Methods
Nie et al. explored the assignment of weights to different graphs for highlighting the significance of key views [21, 22]. These methods follow the technical approach of pursuing consistency for the target graph, which has inspired several variants [51, 52, 23, 53]. Research has shown that establishing a unified graph matrix from different views is useful to integrate comprehensive information [54, 24]. In order to reduce dimensionality, a parameter-free method that allows for structured graph learning was investigated in [55]. In [56], Kang et al. presented a graph learning framework that is able to preserve a complete structure of data. In addition, some research works have tackled other issues. For example, Wen et al. investigated the graph completion issue for incomplete views [57] to explore hidden information in missing views. Xing et al. proposed learning the consensus graph via correntropy to enhance robustness in the presence of missing views. To deal with large-scale datasets, Xia et al. focused on bipartite graph learning to extract fundamental graph information [39]. Wang et al. designed a scalable and flexible anchor graph fusion framework to unify the procedures of anchor graph construction and individual graph fusion, as well provide a more reasonable anchor selection scheme by considering data diversity [58].
@@ III-A Motivation
Observation 1: In order to reflect the similarity among data samples, a common strategy is to learn the adaptive neighbor graph for $\mathbf{A}^{(v)}$ [23, 37]. However, most methods construct the adaptive neighbor graph by Euclidean distance, which may not effectively uncover the inherent structure of high-dimensional data, since using Euclidean distance inevitably suffers from the curse of dimensionality.
Observation 2: We also notice that the majority of graph-based multi-view methods follow the technical approach of learning a consensus graph for the target graph by minimizing the discrepancy between the similarity graph and the target graph. This learning strategy notably emphasizes the importance of consistency among views. However, from another perspective, it neglects the diversity inherent in different views.
Observation 3: Despite a few graph-based works, such as [24], have attempted to consider both consistency and specificity, they model the similarity matrix $\mathbf{A}^{(v)}$ with a consistent part $\mathbf{S}$ and a specific part $\mathbf{S}^{(v)}$, i.e., $\mathbf{A}^{(v)}=\mathbf{S}+\mathbf{S}^{(v)}$. However, this model exists a significant limitation. It does not take into account the impact of noise, which probably leads to the noise being assigned to both the consistent and specific components during the learning process. Ideally, either the consistent or specific component should capture useful information rather than containing noise. Therefore, it is necessary to explicitly distinguish noise from useful components when modeling.
@@ III-B Model Formulation
Stiefel manifold: The Stiefel manifold represents the set of matrices with orthonormal columns. The Stiefel manifold is commonly used in optimization and geometry, particularly in problems involving rotations and orthogonal transformations. Given two $d\times n$ orthonormal matrices $\mathbf{X}$ and $\mathbf{Y}$, the $l_{1}$-norm geodesic distance between them is defined as [59]
$d_{s}(\mathbf{X},\mathbf{Y})=d-\mathrm{tr}(\mathbf{X}\mathbf{Y}^{T}).$ (7)
The above equation characterizes the distance between two matrices $\mathbf{X}$ and $\mathbf{Y}$. However, our goal is to measure the distance between two samples, i.e., two vectors, rather than between matrices. Consequently, it is necessary to adapt Eq. (7) to align with our requirements. A straightforward approach is to set $n=1$, reducing the distance calculation in Eq. (7) to a form suitable for vectors. The resulting expression is as follows:
@@
where $\mathbf{x}$ and $\mathbf{y}$ are two $d\times 1$ vectors. Strictly speaking, since $\mathbf{x}$ and $\mathbf{y}$ are not necessarily unit vectors, this formulation does not rigorously adhere to the definition of the geodesic distance on the Stiefel manifold for the case of $n=1$. Although our distance measure is not a true Stiefel manifold distance, it draws inspiration from Stiefel manifold distance. Therefore, we refer to the distance measure in Eq. (8) as the “pseudo-Stiefel manifold distance”.
@@
Given a set of multi-view data matrices $\{\mathbf{X}^{(v)}\}^{m}_{v=1}$, with $\mathbf{X}^{(v)}\in\mathbb{R}^{d_{v}\times n}$ denoting the $v$-th view data with $d_{v}$ dimensions, we can define a new formulation to learn the neighbor graph $\mathbf{A}^{(v)}$, that is
@@
where $\mathbf{x}^{(v)}_{i}$ represents the $i$-th sample of $\mathbf{X}^{(v)}$, $\alpha$ denotes the positive balance parameter, $\mathbf{A}^{(v)}\in\mathbb{R}^{n\times n}$ stands for the similarity matrix, and $\mathbf{a}^{(v)}_{i}\in\mathbb{R}^{n\times 1}$ is a column vector with its $j$-th element being $a^{(v)}_{ij}$. The main difference between the traditional neighbor graph learning and our method exists in the distance measure. In contrast to the traditional approach that utilizes Euclidean distance $\lVert\mathbf{x}^{(v)}_{i}-\mathbf{x}^{(v)}_{j}\rVert^{2}_{2}$ to measure similarity, our method uses pseudo-Stiefel manifold distance, which more precisely reflects the intrinsic structure for high-dimensional data. In the supplementary material, we present a toy experiment that demonstrates the superiority of using the pseudo-Stiefel manifold distance for high-dimensional data compared to the Euclidean distance.
After learning the neighbor graphs of all the views, we need to fuse these neighbor graphs. As mentioned in Observations 2 and 3 of the motivation subsection, most graph-based multi-view methods directly learn a consensus graph, ignoring the specific information of each view. While a few works consider both consistency and specificity, they fail to distinguish noise from useful information. To address these limitations, we suppose that each neighbor graph $\mathbf{A}^{(v)}$ can be modelled by a consistent component, a view-specific component, and a noise component, i.e.,
@@
For capturing the high-order correlation among different neighbor graphs [60], we formulate the following tensor-based graph fusion framework:
@@
Remark 2: We impose $l_{2,1}$-norm on the noise tensor $\mathcal{E}$ to encourage the columns of each frontal slice $\mathbf{E}^{(v)}$ to be zero, which is expected to minimize the impact of noise [11]. We enforce the consistent tensor $\mathcal{S}_{1}$ with the $t$-SVD based tensor nuclear norm to exploit more shared high-order information across various neighbor graphs [18]. Moreover, we impose $F$-norm on the specific tensor $\mathcal{S}_{2}$ in order to ensure the preservation of view-specific information.
Remark 3: In contrast to the model presented in Observation 3, our model explicitly separates the noise term from the consistent and specific components, which guarantees that the consistent and specific information learned is useful and free from noise. In essence, our modelling technique contributes to noise-free graph fusion. To demonstrate our arguments, we conduct experiments by plotting the probability density function curves for the learned noise-free information and error term under two situations: in the absence of synthetic noise and in the presence of synthetic noise. Additionally, to further validate that noise information is effectively captured in the error component $\mathbf{E}^{(v)}$, we compare two cases for the fused graph matrix: one incorporating only useful information and the other including the error component. The experimental results and their corresponding analysis are provided in the supplemental material.